OpenStudy (anonymous):

(7x-21)/(x^2-7x+12) Please find vertical asymptotes! PLEASE HELP ME!!

OpenStudy (anonymous):

I got x = 3 and x = 4 But whenI graphed it is only x=4 why?

jimthompson5910 (jim_thompson5910):

hint: factor 7x-21 and factor x^2-7x+12

OpenStudy (anonymous):

hint: factor 7x - 21 = ... factor x² -7x+12 = ... Simplify them: ..

OpenStudy (precal):



OpenStudy (precal):

because that (x-3) or x=3 is called a hole

OpenStudy (precal):

see how it cancels out, that tells you it is a hole